# Accelerated creation of NOON states with ultracold atoms via counterdiabatic driving

### v1

@@ Abstract
This project (EOS 40007526) has received funding from the FWO and F.R.S.-FNRS under the Excellence of Science (EOS) programme. S. W. thanks A. Smerzi for useful discussions and acknowledges support by Q-DYNAMO (EU HORIZON-MSCA-2022-SE-01) with project No. 101131418 and by the National Recovery and Resilience Plan (PNRR), Mission 4 Component 2 Investment 1.3 – Call for tender No. 341 of 15/03/2022 of Italian MUR funded by NextGenerationEU, with project No. PE0000023, Concession Decree No. 1564 of 11/10/2022 adopted by MUR, CUP D93C22000940001, Project title “National Quantum Science and Technology Institute“ (NQSTI).
@@
$E\Psi_{0,0,N}=E_{0,0,N}\Psi_{0,0,N}-\mathcal{J}_{N,0}\Psi_{0,1,N-1}$ (7)
This set of equations can be mapped to a set of paths that the system can take to travel through the Hilbert space from $|N,0,0\rangle$ to $|0,0,N\rangle$, as depicted in Fig. 1.
@@
Specific proposal for a realization of a NOON state lattice
@@
At time $t=0$, a quench is to be implemented by which means the on-site energies of the populated lattice sites becomes suddenly increased, thus slightly exceeding the energies on the adjacent sites. As illustrated in Fig. 2(b), this can be achieved by suddenly increasing the amplitude $\eta$ from $-4$ to 00. The resulting (two-period) superlattice configuration can be considered to be the starting point for the NOON state creation process to be implemented. The amplitude $\eta$ has then to be slowly decreased, according to the protocol that we discuss in the main text of the paper. In addition, an artificial gauge field has to be induced along the lattice [1, 2] in order to give rise to the complex hopping matrix element (see Eq. 10 of the manuscript) that is required for the implementation of the counterdiabatic driving protocol. Additional (even time-dependent) adaptations of the Bose-Hubbard parameters can, if needed, be achieved by suitable variations of the global lattice amplitude $V_{0}$ as well as of the analogous amplitudes along the other two axes of the lattice.
As discussed in the main part of the paper, the NOON states are then produced as soon as the populated sites of the lattice have approximately the same energy as their two neighbor sites. As illustrated in Fig. 2(c), this is roughly the case for $\eta=-1$. Further tuning the initially populated (and now depopulated) sites to lower energies then gives rise to a NOON state lattice in which every second site may or may not host $N$ atoms. This can then be verified by quantum gas microscopes [3, 4]. A suitable atom transport protocol [5, 6] can then be used to direct the two components $|N,0\rangle$ and $|0,N\rangle$ of each NOON pair into spatially separate directions, for possible usage in the context of quantum metrological applications.
Note that next-to-nearest neighbor hoppings of atoms, while technically possible, can be safely neglected since they take place on time scales that are much larger than the (already very large) time scales for the creation of the NOON state. The maintaining of nearly perfect homogeneity along the lattice, which is a key requirement for this protocol to work out correctly, is certainly a great challenge but can be, in principle, achieved using flat lattice potentials [7].

### v3

@@ Acknowledgements.
We thank N. Goldman, D. Guéry-Odelin, L. Pezzè and A. Smerzi for inspiring discussions. This project (EOS 40007526) has received funding from the FWO and F.R.S.-FNRS under the Excellence of Science (EOS) programme. S.W. acknowledges support by Q-DYNAMO (EU HORIZON-MSCA-2022-SE-01) with project No. 101131418 and by the National Recovery and Resilience Plan (PNRR), Mission 4 Component 2 Investment 1.3 – Call for tender No. 341 of 15/03/2022 of Italian MUR funded by NextGenerationEU, with project No. PE0000023, Concession Decree No. 1564 of 11/10/2022 adopted by MUR, CUP D93C22000940001, Project title “National Quantum Science and Technology Institute“ (NQSTI).
@@
$E\Psi_{0,0,N}=E_{0,0,N}\Psi_{0,0,N}-\mathcal{J}_{N,0}\Psi_{0,1,N-1},$ (7)
for the components $\Psi_{n_{1},n_{2},n_{3}}$ that are associated with the occupation eigenstates $|n_{1},n_{2},n_{3}\rangle$. This set of equations can be mapped to a set of paths that the system can take to travel through the Hilbert space from $|N,0,0\rangle$ to $|0,0,N\rangle$, as depicted in Fig. 1.
@@
.1 Specific proposal for a realization of a NOON state lattice
@@
At time $t=0$, a quench is to be implemented by which means the on-site energies of the populated lattice sites becomes suddenly increased, thus slightly exceeding the energies on the adjacent sites. As illustrated in Fig. 2(b), this can be achieved by suddenly increasing the amplitude $\eta$ from $-4$ to 00. The resulting (two-period) superlattice configuration can be considered to be the starting point for the NOON state creation process to be implemented. The amplitude $\eta$ has then to be slowly decreased, according to the protocol that we discuss in the main text of the paper. In addition, an artificial gauge field has to be induced along the lattice [1, 2] in order to give rise to the complex hopping matrix element (see Eq. (10) of the manuscript) that is required for the implementation of the counterdiabatic driving protocol. Additional (even time-dependent) adaptations of the Bose-Hubbard parameters can, if needed, be achieved by suitable variations of the global lattice amplitude $V_{0}$ as well as of the analogous amplitudes along the other two axes of the lattice.
As discussed in the main part of the paper, the NOON states are then produced as soon as the populated sites of the lattice have approximately the same energy as their two neighboring sites. As illustrated in Fig. 2(c), this is roughly the case for $\eta=-1$. Further tuning the initially populated (and now depopulated) sites to lower energies then gives rise to a NOON state lattice in which every second site may or may not host $N$ atoms. This can then be verified by quantum gas microscopes [3, 4]. A suitable atom transport protocol [5, 6] can then be used to direct the two components $|N,0\rangle$ and $|0,N\rangle$ of each NOON pair into spatially separate directions, for possible usage in the context of quantum metrological applications.
Note that next-to-nearest neighbor hoppings of atoms, while technically possible, can be safely neglected since they take place on time scales that are much larger than the (already very large) time scales for the creation of the NOON state. Maintaining nearly perfect homogeneity along the lattice, which is a key requirement for this protocol to work out correctly, is certainly a great challenge but can be, in principle, achieved using flat lattice potentials [7].